A circle is divided into 4 arcs of lengths 2,5,6 and x. arc 2 has a measure of 30 degrees. find x
A 8
B 9
C 10
D 11

Answer by lwsshak3(11628)

A circle is divided into 4 arcs of lengths 2,5,6 and x. arc 2 has a measure of 30 degrees. find x
***
arc length=radius*central angle in radians
30º=π/6 radians
2=radius*π/6
radius=12/π
..
Circumference=2π*radius=2π*12/π=24
sum of 4 arcs=24
2+5+6+x=24
13+x=24
x=24-13=11
ans: D 11